OpenStudy (anonymous):

What is the first step in solving the equation −4x − 16 = 20?

OpenStudy (anonymous):

add 16 to both sides

OpenStudy (anonymous):

Oh Ok! Thanks!

OpenStudy (anonymous):

welcome(:

hero (hero):

Add 16 to both sides: -4x = 20 + 16 Combine terms: -4x = 36 Add 4x to both sides; subtract 36 from both sides: 4x = -36 Divide both sides by 4: x = -36/4 x = -9

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

That was what I got! :)